better same worse Republican Democrat none 38 184 12 87 137 21 90 118 Express each of your first five answers as a decimal and round to the nearest 0.001 (in other words, type 0.123, not 12.3% or 0.1234561 What fraction of survey respondents identified themselves as affiliated with neither party? What fraction of survey respondents thought the economy was getting better? What fraction of those affiliated with neither party thought the economy was getting better? Among survey respondents who thought the economy was getting better, what fraction were affillated with neither party? of survey respondents were affillated with neither party who thought the economy was getting better The three ple charts below show the opinions about the economy for each of party
Answer #1

Total number of respondents = 38 + 104 + 44 + 12 + 87 + 137 + 21 + 90 + 118 = 651

Total number of respondents who identified themselves as affiliated with neither party = 21 + 90 + 118= 229

a) The fraction of respondents who identified themselves as affiliated with neither party

229 0.35 18 651

b) Total number of respondents who thought the economy was getting better = 38 + 104 + 44 = 186

The fraction of respondents who thought the economy was getting better

186 0.2857 651

c) Total number of respondents who identified themselves as affiliated with neither party and thought the economy was getting better = 21

The fraction of respondents who identified themselves as affiliated with neither party and thought the economy was getting better

21 651 = 0.0322

d) Among survey respondents who thought the economy was getting better, the fraction of respondents who were affiliated with neither party

21 0.1129 186